OpenStudy (anonymous):

Kyle drives one and two-fifths hours to visit his grandmother when he takes the new expressway, which has made his trip one-third as long as it used to be. How long did his trip take before the expressway was built? one and eleven fifteenths hours seven-fifteenths hours four and one-fifth hours one and one-fifteenth hours

OpenStudy (anonymous):

@campbell_st @Nurali @Luigi0210 can you help meee?

OpenStudy (anonymous):

well...if 1 2/5 hours is 1/3 of 3/3...how do YOU think you should solve the problem?

OpenStudy (anonymous):

multiply 1 2/5 by 3

OpenStudy (anonymous):

which gets you...

OpenStudy (anonymous):

do you know how to multiple mixed numbers by whole numbers?

OpenStudy (anonymous):

no, lol this is a pretest

OpenStudy (anonymous):

I'll explain... change it to an improper fraction (top heavy, which means the numerator is greater than the denomerator) 1 2/5 = 5 x 1 + 2 = 5 + 2 = 7/5 7/5 x 3 = 7/5 x 3/1 = 21/5 21/5 = 4 1/5

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

actually yes!! thankyouu

OpenStudy (anonymous):

when converting mixed numbers to improper fractions you multiply the denomerator by the whole number of the mixed number, add the product to the numerator, and express your solution as anew fraction, with the solution as the numerator over the original denomerator